# Personalized Negative Reservoir for Incremental Learning in Recommender Systems

### v1

@@ Appendix A Case Study Details
Appendix B Loss Ablation Study & Sensitivity Analysis
@@
In this section we conduct a loss ablation study on our proposed objective, as well as sensitivity studies on key components of our method: the choice of specific clustering algorithm, the number of clusters of our clustering algorithm as well as the size of our proposed negative reservoir. The dataset choice for the ablation and sensitivity experiments can be explained by the fact that Netflix, which is our biggest dataset, is prohibitvely big to run many experiments on as it takes over 2 days per experiment (see runtime section), so we opt for Yelp and Taobao14 as they have the most representative number of user and items compared to the average dataset (they are neither the biggest nor the smallest).
@@
Our loss ablation study validates each term in the proposed loss. Concretely, we check the impact of removing each loss term from our overall proposed optimization objective. As we can see in Tables 11 and 12 removing any of our proposed components impacts performance and/or end-to-end trainability.
@@
In the first sensitivity study we investigate if an ad-hoc training scheme where we cluster the items using K-means between epochs can produce similar results to our approach, which uses an end-to-end structural based clustering method (Bo et al., 2020). In Tables 7, 8 we observe that our method performs within 1-3% with either clustering algorithm, thus validating the low sensitivity towards clustering algorithm selection. We obtain similar results when we vary the dataset choice and the number of clusters. The attributed clustering algorithm maintains end-to-end trainability so it converges faster than K-means.
@@
Table 7. Sensitivity to clustering algorithm choice in Yelp. Results demonstrate low sensitivity to specific algorithm choice.
Methods Inc. 1 Inc. 2 Inc.3 Avg. GraphSAIL+SANE (K-means) 0.0877 0.0871 0.0791 0.0846 GraphSAIL+SANE (end-to-end - ours) 0.0939 0.0842 0.0791 0.0857 SGCT+SANE (K-means) 0.0946 0.0858 0.0732 0.0845 SGCT+SANE (end-to-end - ours) 0.0966 0.0877 0.0744 0.0857 LWC-KD+SANE (K-means) 0.0931 0.0853 0.0787 0.0857 LWC-KD+SANE (end-to-end - ours) 0.0970 0.0891 0.0834 0.0898
@@
Table 8. Sensitivity to clustering algo. choice in Taobao14. Results demonstrate low sensitivity to specific algorithm choice.
Methods Inc. 1 Inc. 2 Inc.3 Avg. GraphSAIL+SANE (K-means) 0.0222 0.0139 0.0165 0.0175 GraphSAIL+SANE (end-to-end) 0.0231 0.0131 0.0150 0.0171 SGCT+SANE (K-means) 0.0228 0.0154 0.0153 0.0178 SGCT+SANE (end-to-end) 0.0224 0.0136 0.0173 0.0178 LWC-KD+SANE (K-means) 0.0228 0.0174 0.0154 0.0185 LWC-KD+SANE (end-to-end) 0.0222 0.0188 0.0156 0.0188
@@
In the second sensitivity study, shown in Table 9, we conduct an analysis on the number of clusters used in our method. As we can see, once the number of clusters reaches a sufficient number $\sim 10$, the performance remains stable.
@@
Table 9. Sensitivity analysis for the cluster number K. We present Recall@20 results on the Gowalla and Taobao2015 datasets for SGCT-SANE and LWC-KD-SANE. Results demonstrate low sensitivity to number of clusters.
@@
Table 10. Sensitivity analysis on the size of the user negative reservoir $\lvert\mathcal{Q}\lvert$. We present Recall@20 results on the Gowalla and Taobao2015 datasets for SGCT-SANE and LWC-KD-SANE. Results demonstrate low sensitivity to negative reservoir size.
@@
Table 11. Loss components ablation on Taobao14 and Yelp for SGCT-SANE (Recall@20). We check the impact of removing each loss term from our overall proposed optimization objective. As we can see, removing any of our proposed components impacts performance and/or end-to-end trainability.
Fine Tune Yes 0.0153 0.0661
SGCT Yes 0.0160 0.0668
SGCT-hard-cluster No 0.0178 0.0845
SGCT-SANE (ours) Yes 0.0178 0.0857
Table 12. Loss components ablation on Taobao14 and Yelp for LWCKD-SANE (Recall@20). As we can see, removing any of our proposed components impacts performance and/or end-to-end trainability.
Fine Tune Yes 0.0153 0.0661
LWCKD Yes 0.0176 0.0679
LWCKD-hard-cluster No 0.0185 0.0857
LWCKD-SANE (ours) Yes 0.0188 0.0898
Appendix C Hyperparameter Settings
Our method is implemented in TensorFlow.
Table 13. Hyperparameters of our model on all benchmarks.
@@ Appendix C Hyperparameter Settings
The backbone graph neural network is the MGCCF (Sun et al., 2019) trained using the hyperparameters shown in Table 13. Incremental learning methods are not used during the base block training so the loss during the base block is only $\mathcal{L}_{\textsc{BPR}}$, i.e., (no $\mathcal{L}_{\textsc{KD}}$, $\mathcal{L}_{\textsc{SANE}}$, $\mathcal{L}_{\textsc{KL}}$ components).

### v2

@@ Appendix F Case Study Details
Appendix G Loss Ablation Study & Sensitivity Analysis
G.1 Ablation study
@@
In this section we conduct a loss ablation study on our proposed objective, as well as sensitivity studies on key components of our method: the choice of specific clustering algorithm, the number of clusters of our clustering algorithm as well as the size of our proposed negative reservoir. The dataset choice for the ablation and sensitivity experiments can be explained by the fact that Netflix, which is our biggest dataset, is prohibitively big to run many experiments on as it takes over 2 days per experiment so we opt for Yelp and Taobao14 as they have the most representative number of user and items compared to the average dataset (they are neither the biggest nor the smallest).
@@
Table 5: Loss components ablation on Taobao14 and Yelp for SGCT-SANE (Recall@20). We check the impact of removing each loss term from our overall proposed optimization objective. As we can see, removing any of our proposed components impacts performance and/or end-to-end trainability.
Fine Tune Yes 0.0153 0.0661
SGCT Yes 0.0160 0.0668
SGCT-hard-cluster No 0.0178 0.0845
SGCT-SANE (ours) Yes 0.0178 0.0857
Table 6: Loss components ablation on Taobao14 and Yelp for LWCKD-SANE (Recall@20). As we can see, removing any of our proposed components impacts performance and/or end-to-end trainability.
Fine Tune Yes 0.0153 0.0661
LWCKD Yes 0.0176 0.0679
LWCKD-hard-cluster No 0.0185 0.0857
LWCKD-SANE (ours) Yes 0.0188 0.0898
Our loss ablation study validates each term in the proposed loss. Concretely, we check the impact of removing each loss term from our overall proposed optimization objective. As we can see in Tables 5 and 6 removing any of our proposed components impacts performance and/or end-to-end trainability.
G.2 Clustering Algorithm Sensitivity
@@
In this sensitivity study we check if an ad-hoc training scheme where we cluster the items using K-means between epochs can produce similar results to our approach, which uses an end-to-end structural based clustering method (Bo et al., 2020). In Table 8 observe that our method performs within 1-3% with either clustering algorithm, thus validating the low sensitivity towards clustering algorithm selection. We obtain similar results when we vary the dataset choice and the number of clusters. The attributed clustering algorithm maintains end-to-end trainability so it converges faster than K-means.
@@
Table 7: Sensitivity to clustering algorithm choice in Yelp. Results demonstrate low sensitivity to specific algorithm choice.
Methods Inc. 1 Inc. 2 Inc.3 Avg.
GraphSAIL+SANE (K-means) 0.0877 0.0871 0.0791 0.0846
GraphSAIL+SANE (end-to-end - ours) 0.0939 0.0842 0.0791 0.0857
SGCT+SANE (K-means) 0.0946 0.0858 0.0732 0.0845
SGCT+SANE (end-to-end - ours) 0.0966 0.0877 0.0744 0.0857
LWC-KD+SANE (K-means) 0.0931 0.0853 0.0787 0.0857
LWC-KD+SANE (end-to-end - ours) 0.0924 0.0855 0.0798 0.0859
@@
Table 8: Sensitivity to clustering algo. choice in Taobao14. Results demonstrate low sensitivity to specific algorithm choice.
Methods Inc. 1 Inc. 2 Inc.3 Avg.
GraphSAIL+SANE (K-means) 0.0222 0.0139 0.0165 0.0175
GraphSAIL+SANE (end-to-end) 0.0231 0.0131 0.0150 0.0171
SGCT+SANE (K-means) 0.0228 0.0154 0.0153 0.0178
SGCT+SANE (end-to-end) 0.0224 0.0136 0.0173 0.0178
LWC-KD+SANE (K-means) 0.0228 0.0174 0.0154 0.0185
LWC-KD+SANE (end-to-end) 0.0222 0.0188 0.0156 0.0188
@@
In the second sensitivity study, shown in Table 9, we conduct an analysis on the number of clusters used in our method. As we can see, once the number of clusters reaches a sufficient number ($\sim 10$), the performance remains stable.
@@
Table 9: Sensitivity analysis for the cluster number K. We present Recall@20 results on the Gowalla and Taobao2015 datasets for SGCT-SANE and LWC-KD-SANE. Results demonstrate low sensitivity to number of clusters.
@@
Table 10: Sensitivity analysis on the size of the user negative reservoir $\lvert\mathcal{Q}\lvert$. We present Recall@20 results on the Gowalla and Taobao2015 datasets for SGCT-SANE and LWC-KD-SANE. Results demonstrate low sensitivity to negative reservoir size.
@@
Appendix H Hyperparameter Settings
Our method is implemented in TensorFlow. The backbone graph neural network is the MGCCF (Sun et al., 2019) trained using the hyperparameters shown in Table 11. Incremental learning methods are not used during the base block training so the loss during the base block is only $\mathcal{L}_{\textsc{BPR}}$, i.e., (no $\mathcal{L}_{\textsc{KD}}$, $\mathcal{L}_{\textsc{SANE}}$, $\mathcal{L}_{\textsc{KL}}$ components).
Table 11: Hyperparameters of our model on all benchmarks.
@@ Appendix H Hyperparameter Settings
Appendix I Experimental Results per Incremental Training Block
In this section we provide a more detailed version of the results from Tab. 1 with a breakdown of the performance for each incremental block.
Table 12: Comparison (Recall@20) of baselines and 3 recent knowledge distillation algorithms with our SANE reservoir, our methods are accompanied by a star (*). Best performers are in bold, second best are underlined. Blue cell colors indicate improvement of a particular method when our reservoir is introduced.
Yelp Netflix
Model Inc 1 Inc 2 Inc 3 Average $\Delta$Inc 1 Inc 2 Inc 3 Average $\Delta$%
Fine Tune 0.0705 0.0638 0.0640 0.0661 0.0 0.1092 0.1041 0.0977 0.1036 0.0
LSP_s 0.0722 0.0661 0.0644 0.0676 2.3 0.1173 0.1136 0.1076 0.1128 8.7
Uniform 0.0718 0.0635 0.0610 0.0654 -1.1 0.1018 0.1055 0.0800 0.0957 -7.6
Inv_degree 0.0727 0.0699 0.0605 0.0677 2.4 0.1000 0.1050 0.0820 0.0957 -7.6
GraphSAIL 0.0674 0.0617 0.0625 0.0639 -3.3 0.1163 0.1023 0.0968 0.1051 1.4
GraphSAIL-SANE* 0.0939 0.0842 0.0791 0.0857 29.7 0.1153 0.1091 0.01014 0.1086 4.8
SGCT 0.0740 0.0656 0.0608 0.0668 1.06 0.1166 0.1161 0.1077 0.1135 9.6
SGCT-SANE* 0.0966 0.0877 0.0744 0.0862 30.4 0.1182 0.1164 0.1120 0.1155 11.5
LWC-KD 0.0739 0.0661 0.0637 0.0679 2.7 0.1185 0.1170 0.1071 0.1142 10.2
LWC-KD-SANE* 0.0970 0.0891 0.0834 0.0898 35.9 0.1192 0.1196 0.1157 0.1182 14.1
Taobao14 Taobao15
Model Inc 1 Inc 2 Inc 3 Average $\Delta$Inc 1 Inc 2 Inc 3 Average $\Delta$%
Fine Tune 0.0208 0.0112 0.0138 0.0153 0.0 0.0933 0.0952 0.0965 0.0950 0.0
LSP_s 0.0213 0.0106 0.0138 0.0152 -0.7 0.0993 0.0952 0.0957 0.0968 1.9
Uniform 0.0195 0.0127 0.0148 0.0157 2.6 0.0988 0.0954 0.1004 0.0982 3.4
Inv_degree 0.0228 0.0140 0.0159 0.0175 14.6 0.0991 0.0977 0.1000 0.0989 4.2
GraphSAIL 0.0222 0.0105 0.0139 0.0155 1.31 0.0959 0.0959 0.0972 0.0963 1.4
GraphSAIL-SANE* 0.0231 0.0131 0.0150 0.0171 11.8 0.1114 0.1087 0.1121 0.1107 16.5
SGCT 0.0240 0.0092 0.0148 0.0160 1.74 0.1030 0.0983 0.0984 0.0999 5.2
SGCT-SANE* 0.0224 0.0136 0.0173 0.0178 16.3 0.1117 0.1129 0.1138 0.1128 18.7
LWC-KD 0.0254 0.0119 0.0156 0.0176 15.3 0.1039 0.1022 0.1029 0.1030 8.4
LWC-KD-SANE* 0.0222 0.0188 0.0155 0.0188 22.9 0.1106 0.1108 0.1128 0.1114 17.2
Gowalla-20 MovieLens10M
Model Inc 1 Inc 2 Inc 3 Average $\Delta$Inc 1 Inc 2 Inc 3 Average $\Delta$%
Fine Tune 0.1412 0.1637 0.2065 0.1705 0.0 0.0923 0.0904 0.0957 0.0918 0.0
LSP_s 0.1512 0.1741 0.2097 0.1783 4.57 0.0912 0.0923 0.0941 0.0925 0.76
Uniform 0.1480 0.1653 0.2051 0.1728 1.34 0.0874 0.0879 0.0902 0.0885 -3.59
Inv_degree 0.1483 0.1680 0.2001 0.1738 1.93 0.0906 0.0911 0.0928 0.0915 -0.32
GraphSAIL 0.1529 0.1823 0.2195 0.1849 8.44 0.0867 0.0887 0.0917 0.0890 -3.05
GraphSAIL-SANE* 0.1646 0.1907 0.2221 0.1925 12.9 0.0887 0.0915 0.0961 0.0921 0.33
SGCT 0.1588 0.1815 0.2207 0.1870 9.68 0.0894 0.0902 0.0960 0.0919 0.11
SGCT-SANE* 0.1611 0.1843 0.2237 0.1897 11.3 0.0927 0.0932 0.0980 0.0946 3.05
LWC-KD 0.1639 0.1921 0.2368 0.1977 15.9 0.0902 0.0910 0.0973 0.0928 1.09
LWC-KD-SANE* 0.1698 0.1835 0.2173 0.1881 10.3 0.0928 0.0953 0.1012 0.0964 5.01
Appendix J Dataset Preprocessing - Data Block Definition
For all data sets we follow the setup from literature (Xu et al., 2020; Ahrabian et al., 2021). We divide the data in chronological order into a base block and incremental blocks. The base block contains 60% of the data. The remaining 40% of the data is evenly separated into four incremental blocks. Note that during training the blocks can be further subdivided to allow for a validation set. In our setting for each block we reserve 5% of the data as a validation dataset following from the baselines (Xu et al., 2020; Ahrabian et al., 2021). Note that this does not induce data leakage as all splits are in chronological order. For a visualization see Fig. 5.
Figure 5: Data is separated to a base block with 60% of the data and 4 incremental blocks, each with 10% of the remaining data.
Appendix K Practical Recommendation for Model Deployment
Tuning the framework hyperparameters can be achieved in similar fashion to the setup in Fig. 5. Here, we again split the data into temporally sorted blocks. The base block contains 60% of the data. The remaining 40% of the data is evenly separated into four incremental blocks. During training, when block $t$ is used as the training set, the first half of block $t+1$ is used as the validation set and the second half is used as the testing set (Figure 6). Note this setup prevents data leakage as the test data are all from interactions that occur after the validation set. For example, if a block contains a 10 days worth of data, the validation dataset would contain days 1-5 and the test set days 6-10.
Figure 6: Incremental learning hyperparameter tuning setup. Note this setup prevents data leakage as the test data are all from interactions that occur after the validation data. For example, if a block contains a 10 days worth of data, the validation dataset would contain days 1-5 and the test set days 6-10.
To deploy GraphSANE on top of an existing recommendation system, we recommend the following steps:
  1. 1.
    Accumulate data equal to 10 times the incremental block size. For example, if the model is expected to be updated daily we recommend accumulating 10 days worth of data for hyperparameter tuning.
  2. 2.
    Split the data chronologically in blocks as shown in Fig. 6. The reason that we put validation sets in the data immediately following the block is the we wish to create a back-testing style setup where the hyperpameters are tuned based on the typical distribution shift between the train and test steps. Note that since the blocks are chronologically ordered this does not leak any interaction from the test set to the validation set. This placement of the validation set is a practice adopted form the literature (Xu et al., 2020; Ahrabian et al., 2021).
  3. 3.
    Train a base GNN model on the base block data. Note that any base backbone recommendation system model, such as MGCCF (Sun et al., 2019) or LightGCN (He et al., 2020b) has its own hyperparameters which can be tuned on the validation portion of the base block. These base GNN architectures have their own hyperparameters that are not related to our incremental learning framework so we omit further discussion on how to tune the base models and focus on how to shift hyperparameters related to our method.
  4. 4.
    Tune $\beta$ parameter by checking the training loss of the first incremental block. It suffices to select $\beta$ values such that at the start of training the KL loss component of the loss is in the same scale as the triplet loss component. Note that the source code we provide tracks these components so this process is easy to perform.
  5. 5.
    For the number of clusters $K$ and the size of the reservoir $|Q|$ we recommend running a grid search. Note that since these are two hyperparameters we may run a search over the values we used in the sensitivity tests we run in Tables 9 and 10. Note even though we run a grid search in our experiments, the parameter tuning process could be further optimized using Bayesian optimization hyperparameter tuning tools such as Optuna  [^2] https://optuna.org/.